OpenStudy (anonymous):

Solve by substitution 5x-21=-y 2x-y=0

jhonyy9 (jhonyy9):

y=2x

OpenStudy (amistre64):

what does -y equal?

jhonyy9 (jhonyy9):

than 5x+2x=21 7x=21 x=21/7 x=3

OpenStudy (anonymous):

Thats the problem you are suppose to solve by substitution

OpenStudy (amistre64):

do they tell you what -y equals or do you just have to guess at it?

jhonyy9 (jhonyy9):

so than y= 2*3 =6 y=6

OpenStudy (anonymous):

it doesnt say

OpenStudy (amistre64):

perhaps your not understanding the question im asking :) it isnt about what solid value you can attribute to -y like saying -y=28 or some such; but its a related value

OpenStudy (amistre64):

does it give you some sort of indication as to what -y COULD be?

OpenStudy (anonymous):

thats the exact from my paper what i type

OpenStudy (amistre64):

im sure it is ...

OpenStudy (amistre64):

5x-21= -y would you agree from this here that -y equals some relation to 5x-21 ?

OpenStudy (anonymous):

x = 21/5, y = 0

OpenStudy (amistre64):

this is key to understanding how substitution works to solve these

OpenStudy (amistre64):

it tells you in the problem that "-y" is equal to some relation such that "-y" equals "5x-21".

OpenStudy (amistre64):

and since -y = 5x-21; wherever we see a "-y" we can replace it by "5x-21" the second equation gives us this situation: 2x -y = 0 ; we can substitute in the "value" of -y in here since it is equal to it to begin with 2x + (5x-21) = 0

OpenStudy (anonymous):

ok heres another problem solve by substitution 15x-20y=-1 5y=-1+10x

OpenStudy (amistre64):

how are we to solve it?

OpenStudy (anonymous):

by substitution

OpenStudy (amistre64):

by substitution doesnt mean that you get someone else to solve it for you :)

OpenStudy (amistre64):

what value can we relate to x or y ?

OpenStudy (anonymous):

duh im here for help not sarcasm

OpenStudy (amistre64):

then take the help ....

OpenStudy (anonymous):

i dont need yours goodbye

OpenStudy (amistre64):

oh but you do, you do ... by this type of problem, we should e familiar with how to manipulate and equation to solve it for a particular part thru algebraic means

OpenStudy (amistre64):

if we cant yet solve for either an x or a y in terms of the information given; then there is no real hope of getting to a solution

OpenStudy (anonymous):

toodles someone will have a better attitude then yours

OpenStudy (amistre64):

by chance they might; or they might be satisfied with simply touting off an answer and leaving you in your ignorance.

OpenStudy (amistre64):

i would rather attempt to teach you how to get a solution for yourself and empower you for the future

OpenStudy (anonymous):

i am going to report you thanks for the comments.

OpenStudy (amistre64):

and your report will be sent to me in an email and I will read it and make a choice on what action I should take as a result ... that is what a moderator does

OpenStudy (amistre64):

in the mean time, you will continue to sidestep the real issue here and take offense to someone actually taking the time to try to enlighten you

OpenStudy (anonymous):

We'll see about that

OpenStudy (amistre64):

Now given your equations: 15x-20y=-1 5y=-1+10x do you recall any means whereby we can single out an x or a y that we can use to substitute into the other with?

OpenStudy (amistre64):

given that i dont have the time needed to sit here and try to help you; i have to run to class for a few hours. Maybe someone will come along and give you the guidance you are seeking. Good luck with it :)